OpenStudy (anonymous):

g(x)= -0.5x^2+3x+4 when g(x) =-5. trying to find x-intercepts.

OpenStudy (kira_yamato):

No real values of solutions (x-intercepts)...

OpenStudy (anonymous):

Not an option. =\ It's online homework.

OpenStudy (anonymous):

-5 = -0.5x^2 +3x + 4 -0.5x^2 +3x + 9 = 0 -x^2 + 6x + 18 = 0 x1,2 = (-6 +/- (36 + 72)^0.5) / -2 x1,2 = 3 +/- (27)^0.5 x1,2 = 3 +/- 3*(3^0.5)

OpenStudy (anonymous):

so x=6 or x=0?

OpenStudy (anonymous):

no 3 + 3* (3^0.5) = 8.196.... 3 - 3 * (3^0.5) = -2.196..

OpenStudy (anonymous):

instead of 3 * (3^0.5) you can say 27^0.5 hope it is better for you now

OpenStudy (anonymous):

Thank you so much!!